# LitE-SNN: Designing Lightweight and Efficient Spiking Neural Network through Spatial-Temporal Compressive Network Search and Joint Optimization

### v1

@@ 3.4 Loss Function and Joint Optimization Algorithm
The objective of LitESNN is to identify SNNs that are both lightweight and efficient, achieving high performance while minimizing memory and computation costs. This can naturally be formulated as a multi-objective optimization problem. We formulate the following loss function:
@@
where the coefficient $\lambda_{1}$ and $\lambda_{2}$ serve as regularization parameters that control the trade-off among $Loss_{CE}$, $Loss_{MEM}$ and $Loss_{COMP}$. The first term $Loss_{CE}$ represents the cross-entropy (CE) loss, responsible for evaluating the predictive accuracy of the SNN. Following [?], the original definition of $Loss_{CE}$ is:
@@
The second term $Loss_{MEM}$ quantify the memory usage of the network, which is measured by the model size:
@@
$b_{w}=\sum_{j}^{len(B)}\frac{exp\big{(}\beta^{(j)}\big{)}}{\sum_{j}^{m}exp\big{(}\beta^{(j)}\big{)}}\cdot q^{(j)}$ (19)
where $q^{j}$ denotes the $j$-th bitwidth candidate. The third term $Loss_{COMP}$ assesses the computational cost, which is affected not only by the number of synaptic operations (SynOps) but also by the operational precision [?]. Therefore, we define this term as bit-SynOps, reflecting the computational complexity of the model. The corresponding equation is:
@@
The joint optimization comprises two steps for each iteration. First, we conduct the first forward pass and update the weights, which constitute the foundational components of the network. Then, a second forward pass is conducted using the updated weights and we simultaneously update $\beta$ and $mask$ within our CompConv, as well as the $\psi$ and architecture parameters $\alpha$. Both two steps apply the loss function in Equation (15). Our joint optimization ensures cohesive and efficient optimization, avoiding local optimal and suboptimal global solutions resulting from multiple single-objective optimizations [?].
After search, we decode the cell structure (in Figure 1(a)) by retaining the two strongest incoming edges for each node and select the node operation (in Figure 1(b)) with the strongest edge, where the strength of edge is determined by $\alpha$. The bit-width (in Figure 2(b)) chosen for each cell is decoded by the branch with the highest weight $\beta$. We also select the number of timesteps with the highest weight $\psi$ (in Figure 2(d)). During the retraining, we use the auxiliary loss as in [?]. The weights and masks are retrained to adapt to single-branch network. We adopt the surrogate gradient to direct train the SNN and use the Dspike function [?] to approximate the derivative of spike activity.
@@ 4.1 Experimental Settings
Datasets. We evaluate our proposed LitESNN on two image datasets: CIFAR-10 [?], CIFAR-100 [?]. We also choose an audio dataset Google Speech Command (GSC) [?] which represents the keyword-spotting task commonly deployed on edge devices for quick responses.
CIFAR-10 has 60,000 images of 10 classes with a size of 32$\times$32. Among them, there are 50,000 training images and 10,000 testing images. CIFAR-100 has the same configurations as CIFAR-10, except it contains 100 classes. We split the dataset into 9,000 training samples and 1,000 test samples. The pre-processing method we use on these two datasets is the same as [?].
GSC has 35 words spoken by 2,618 speakers. Following the pre-processing method used in related studies [?], we split the dataset into 12 classes, which include 10 keywords and two additional classes. We apply Mel Frequency Cepstrum Coefficient (MFCC) to extract acoustic features. The sampling frequency is $16$kHz, the frame length and shift are set to $30$ms and $10$ms, and the filter channel is defined as $40$.
Hyperparameters. We implement our LitESNN using Pytorch on NVIDIA GeForce RTX 3080 (10G) GPUs. During the architecture search, we conduct $50$ epochs with a batch size of $30$. We use the SGD optimizer with momentum $0.9$ and a learning rate of $0.025$ to update network weights $w$ and employ the Adam optimizer with a learning rate of $3e^{-4}$ to update $\beta$, $mask$, and $\alpha$. During the retraining, we train for $200$ epochs with a batch size of $50$, using the SGD optimizer with momentum $0.9$ and a cosine learning rate of $0.025$. The neuron parameters $\tau_{decay}$ and $V_{reset}$ are set to $0.2$ and 00 respectively. The threshold $V_{th}$ is initialized to $0.5$. The temperature parameter in Dspike function is set to $3$ and the settings of other parameters follow [?].
@@ 4.2 Comparison with Existing SNNs
Table 2111The results of comparison SNNs are from original papers, or (if not provided) from our experiments using publicly available code from the paper. If the original paper reported multiple results, we list two results with the highest accuracy in the table. presents a comprehensive comparison between our proposed LitESNNs and other state-of-the-art lightweight SNNs on CIFAR10 and CIFAR-100 datasets. Our large-scale LitESNN achieves state-of-the-art accuracy on CIFAR-10 and CIFAR-100 datasets while maintaining a remarkably compact model size of 3.60MB and 3.62MB and relatively fewer BitSynOps of 2863M and 3590M. We can see that, this compressed SNN achieves superior accuracy compared with the baseline model without compression [?]. This result reveals the potential of compression to remove noisy and irrelevant connections to enhance the ability to capture valuable information. Our small-scale LitESNN has a model size of only 0.55MB and Bit-SynOps of 298M, smaller than all other models except the model size of [?] on CIFAR-10 dataset. Nevertheless, it achieves a competitive accuracy of 91.98%, outperforming [?] by a big margin of 8.63%. Notice that [?], [?] and [?] also maintain a good performance with relatively small model size and Bit-SynOps. This is because they employ effective pruning methods to reduce the number of parameters. Despite this, under similar model sizes or Bit-SynOps, our LitESNNs still achieve higher accuracy. For example, on CIFAR-10 dataset, our medium-scale LitESNN achieves 94.52% accuracy with a 1.23MB model, surpassing [?] by 1% while being 1.6MB smaller. The results of our timestep search are also shown in Table 2. From large-scale to small-scale LitESNN, we observe that the optimal number of timesteps decreases from 6 to 4. This is because the increased $\lambda_{1}$ and $\lambda_{2}$ values in the search impose stricter resource limits. Consequently, the network tends to select a smaller number of timesteps to minimize calculations.
@@ 4.3 Effectiveness of $\lambda_{1}$ and $\lambda_{2}$
In this section, we vary $\lambda_{1}$ and $\lambda_{2}$ in Equation (15) and validate its effect on the trade-off between accuracy and resource costs in the LitESNN on CIFAR-10 dataset. The other parameters maintain the same setting as our small-scale LitESNN.
@@ 4.4 Comparison with Sequential Optimization
LitESNN (large) 95.60 3.60 2863 5.01
LitESNN (medium) 94.52 1.23 913 4.99
LitESNN (small) 91.98 0.55 298 4.48
@@
In this section, we conduct a comparative analysis between our LitESNNs and the SNNs optimized using the sequential scheme on CIFAR-10 dataset to validate the effectiveness our joint optimization scheme. The comparison models are derived by sequentially performing neural architecture search (A), pruning (P), quantization (Q) and timestep selection (T). The methods used in each step follow the corresponding scales of LitESNNs. We also conduct a comparison of the total time required to design the network from scratch (including search and retraining).
As shown in Table 4, the three versions of LitESNN consistently outperform models that are derived from sequential optimization, achieving higher accuracy with comparable model sizes and Bit-SynOps. Since each step in sequential models applies the same methods as LitESNNs, their inferior accuracy and longer design time demonstrate the effectiveness of our joint optimization. It is worth noting that the small-scale sequential model suffers a substantial accuracy loss. This can be attributed to its high sparsity of 90%, making it sensitive to parameter adjustments. Small changes in parameter values arising from subsequent quantization may have a large impact on the network. Whereas our joint optimization takes into account the intertwined effects of various compressive factors, thereby effectively mitigating accuracy degradation.
@@ 4.5 Energy Consumption
In this section, we estimate the theoretical energy consumption of our LitESNN using the common approach in the neuromorphic community [??]. The energy is measured in 45nm CMOS technology [?], where the addition operation costs 0.9 pJ energy and the multiplication operation costs 3.7 pJ. These energy values are based on 32-bit floating-point operations. Since our model uses quantization to a smaller bit width, the reported energy values for our model are overestimates compared to those of other models.
Table 5 presents the energy consumption of our LitESNNs and corresponding ANN counterpart with the same network architecture. Our LitESNNs with three scales respectively cost 0.68 mJ, 0.26 mJ and 0.13 mJ for a single forward and consumes over 6$\times$ lower energy compared with ANN. Compared with our baseline model SpikeDHS, LitESNNs achieve an energy reduction of more than 0.58mJ. We also compare our results with AutoSNN, which focuses on designing energy-efficient SNNs. Its energy consumption is comparable with our large-scale LitESNN, larger than our small and medium-scale LitESNNs. It’s essential to acknowledge that energy consumption is not only from computation but also from memory access, a topic beyond the scope of our study as it is more about hardware design. Nevertheless, it’s worth noting that our model’s smaller size and lower bit-width can contribute to reducing access-related energy consumption.
@@
AutoSNN [?] 585.60 28.31 0.63
SpikeDHS [?] 1305.65 23.89 1.26
LitESNN (large) 705.01 11.94 0.68
LitESNN (medium) 271.41 4.78 0.26
LitESNN (small) 129.74 2.39 0.13
@@ 5 Conclusion
In this paper, we propose the LitESNN to integrate spatio-temporal compression into the automatic network design. We present the CompConv that expands the search space to support pruning and mixed-precision quantization. Meanwhile, we propose the compressive timestep search to identify the optimal number of timesteps to process SNNs. We also formulate a joint optimization algorithm to learn these compression strategies and architecture parameters. Experimental results show that our LitESNN exhibits competitive accuracy, compact model size, and fewer computation costs, making it an attractive choice for deployment on resource-constrained edge devices and opening new possibilities for the practical implementation of SNNs in real-world applications.

### v2

@@ 3.4 Loss Function and Joint Optimization Algorithm
The objective of LitE-SNN is to identify SNNs that are both lightweight and efficient, achieving high performance while minimizing memory and computation costs. This can naturally be formulated as a multi-objective optimization problem. We formulate the following loss function:
@@
where the coefficients $\lambda_{1}$ and $\lambda_{2}$ serve as regularization parameters that control the trade-off among $Loss_{CE}$, $Loss_{MEM}$ and $Loss_{COMP}$. The first term $Loss_{CE}$ represents the cross-entropy (CE) loss, responsible for evaluating the predictive accuracy of the SNN. Following Che et al. (2022), the original definition of $Loss_{CE}$ is:
@@
The second term $Loss_{MEM}$ quantifies the memory usage of the network, which is measured by the model size:
@@
$b_{w}=\sum_{j}^{len(B)}\frac{exp\big{(}\beta^{(j)}\big{)}}{\sum_{j}^{len(B)}exp\big{(}\beta^{(j)}\big{)}}\cdot q^{(j)}$ (19)
where $q^{j}$ denotes the $j$-th bitwidth candidate. The third term $Loss_{COMP}$ assesses the computational cost, which is affected not only by the SynOps but also by the operational precision Cai and Vasconcelos (2020). Therefore, we define this term as bit-SynOps, reflecting the computational complexity of the model. The corresponding equation is:
@@
The joint optimization comprises two steps for each iteration. First, we conduct the first forward pass and update the weights, which constitute the foundational components of the network. Then, a second forward pass is conducted using the updated weights and we simultaneously update $\beta$ and $mask$ within our CompConv, as well as the $\psi$ and architecture parameters $\alpha$. Both two steps apply the loss function in Equation (15). Our joint optimization ensures cohesive and efficient optimization, avoiding local optimal and suboptimal global solutions resulting from multiple single-objective optimizations Wang et al. (2020).
After search, we decode the cell structure (in Figure 1(a)) by retaining the two strongest incoming edges for each node and select the node operation (in Figure 1(b)) with the strongest edge, where the strength of edge is determined by $\alpha$. The bit-width (in Figure 2(b)) chosen for each cell is decoded by the branch with the highest weight $\beta$. We also select the number of timesteps with the highest weight $\psi$ (in Figure 2(d)). During the retraining, we use the auxiliary loss as in Che et al. (2022). The weights and masks are retrained to adapt to single-branch network. We adopt the surrogate gradient to direct train the SNN and use the Dspike function Li et al. (2021) to approximate the derivative of spike activity.
@@ 4 Experiments
This section compares our model’s performance and efficiency with current benchmarks and optimization methods and evaluates the effectiveness of parameters $\lambda_{1}$ and $\lambda_{2}$.
@@ 4.1 Experimental Settings
Datasets. We evaluate our proposed LitE-SNN on two image datasets: CIFAR-10 Krizhevsky et al. (2009), CIFAR-100 Krizhevsky et al. (2009). We also choose an audio dataset Google Speech Command (GSC) Warden (2018) which represents the keyword-spotting task commonly deployed on edge devices for quick responses.
CIFAR-10 has 60,000 images of 10 classes with a size of 32$\times$32. Among them, there are 50,000 training images and 10,000 testing images. CIFAR-100 has the same configurations as CIFAR-10, except it contains 100 classes. We split the dataset into 9,000 training samples and 1,000 test samples. The pre-processing method we use on these two datasets is the same as Che et al. (2022).
GSC has 35 words spoken by 2,618 speakers. Following the pre-processing method used in Yang et al. (2022), we split the dataset into 12 classes, which include 10 keywords and two additional classes. We apply Mel Frequency Cepstrum Coefficient (MFCC) to extract acoustic features. The sampling frequency is $16$kHz, the frame length and shift are set to $30$ms and $10$ms, and the filter channel is defined as $40$.
Hyperparameters. We implement our LitE-SNN using Pytorch on NVIDIA GeForce RTX 3080 (10G) GPUs. During the architecture search, we conduct $50$ epochs with a batch size of $30$. We use the SGD optimizer with momentum $0.9$ and a learning rate of $0.025$ to update network weights $w$ and employ the Adam optimizer with a learning rate of $3e^{-4}$ to update $\beta$, $mask$, and $\alpha$. During the retraining, we train for $200$ epochs with a batch size of $50$, using the SGD optimizer with momentum $0.9$ and a cosine learning rate of $0.025$. The neuron parameters $\tau_{decay}$ and $V_{reset}$ are set to $0.2$ and 00 respectively. The threshold $V_{th}$ is initialized to $0.5$. The temperature parameter in Dspike function is set to $3$ and the settings of other parameters follow Che et al. (2022).
@@ 4.2 Comparison with Existing SNNs
Table 2111The results of comparison SNNs are from original papers, or (if not provided) from our experiments using publicly available code from the paper. If the original paper reported multiple results, we list two results with the highest accuracy in the table. presents a comprehensive comparison between our proposed LitE-SNNs and other state-of-the-art lightweight SNNs on CIFAR10 and CIFAR-100 datasets. Our large-scale LitE-SNN achieves state-of-the-art accuracy on CIFAR-10 and CIFAR-100 datasets while maintaining a remarkably compact model size of 3.60MB and 3.62MB and relatively fewer BitSynOps of 2863M and 3590M. We can see that, this compressed SNN achieves superior accuracy compared with the baseline model without compression Che et al. (2022). This result reveals the potential of compression to remove noisy and irrelevant connections to enhance the ability to capture valuable information. Our small-scale LitE-SNN has a model size of only 0.55MB and Bit-SynOps of 298M, smaller than all other models except the model size of Rueckauer et al. (2017) on CIFAR-10 dataset. Nevertheless, it achieves a competitive accuracy of 91.98%, outperforming Rueckauer et al. (2017) by a big margin of 8.63%. Notice that Chen et al. (2021), Kim et al. (2022b) and Chen et al. (2022) also maintain a good performance with relatively small model size and Bit-SynOps. This is because they employ effective pruning methods to reduce the number of parameters. Despite this, under similar model sizes or Bit-SynOps, our LitE-SNNs still achieve higher accuracy. For example, on CIFAR-10 dataset, our medium-scale LitE-SNN achieves 94.52% accuracy with a 1.23MB model, surpassing Kim et al. (2022b) by 1% while being 1.6MB smaller. The results of our timestep search are also shown in Table 2. From large-scale to small-scale LitE-SNN, we observe that the optimal number of timesteps decreases from 6 to 4. This is because the increased $\lambda_{1}$ and $\lambda_{2}$ values in the search impose stricter resource limits. Consequently, the network tends to select a smaller number of timesteps to minimize calculations.
@@ 4.3 Effectiveness of $\lambda_{1}$ and $\lambda_{2}$
We vary $\lambda_{1}$ and $\lambda_{2}$ in Equation (15) and validate its effect on the trade-off between accuracy and resource costs in the LitE-SNN on CIFAR-10 dataset. The other parameters maintain the same setting as our small-scale LitE-SNN.
@@ 4.4 Comparison with Sequential Optimization
LitE-SNN (large) 95.60 3.60 2863 5.01
LitE-SNN (medium) 94.52 1.23 913 4.99
LitE-SNN (small) 91.98 0.55 298 4.48
@@
In this section, we conduct a comparative analysis between our LitE-SNNs and the SNNs optimized using the sequential scheme on CIFAR-10 dataset to validate the effectiveness of our joint optimization scheme. The comparison models are derived by sequentially performing neural architecture search (A), pruning (P), quantization (Q) and timestep selection (T). The methods used in each step follow the corresponding scales of LitE-SNNs.
As shown in Table 4, the three versions of LitE-SNN consistently outperform models that are derived from sequential optimization, achieving higher accuracy with comparable model sizes and Bit-SynOps. Since each step in sequential models applies the same methods as LitE-SNNs, their inferior accuracy and longer design time demonstrate the effectiveness of our joint optimization. It is worth noting that the small-scale sequential model suffers a substantial accuracy loss. This can be attributed to its high sparsity of 90%, making it sensitive to parameter adjustments. Small changes in parameter values due to subsequent quantization can significantly impact the network. Whereas our joint optimization considers the intertwined effects of various compressive factors, thereby effectively mitigating accuracy degradation.
@@ 4.5 Energy Consumption
In this section, we estimate the theoretical energy consumption of our LitE-SNN using the common approach in the neuromorphic community Rathi and Roy (2021); Li et al. (2021). The energy is measured in 45nm CMOS technology Horowitz (2014), where the addition operation and multiplication operation cost 0.9 pJ and 3.7 pJ energy respectively. These energy values are based on 32-bit floating-point operations. Since our model uses quantization to a smaller bit width, the reported energy values for our model are overestimates compared to those of other models.
Table 5 presents the energy consumption of our LitE-SNNs and corresponding ANN counterparts with the same network architecture. Our LitE-SNNs with three scales respectively consume over 6$\times$ lower energy compared with ANN. Compared with our baseline model SpikeDHS, LitE-SNNs achieve an energy reduction of more than 0.58mJ. We also compare our results with AutoSNN, which focuses on designing energy-efficient SNNs. Its energy consumption is comparable with our large-scale LitE-SNN, larger than our small and medium-scale LitE-SNNs. We acknowledge that energy consumption is not only from computation but also from memory access, a topic beyond the scope of our study as it is more about hardware design. Nevertheless, it’s worth noting that our model’s smaller size and lower bit-width can contribute to reducing access-related energy consumption.
@@
AutoSNN Na et al. (2022) 585.60 28.31 0.63
SpikeDHS Che et al. (2022) 1305.65 23.89 1.26
LitE-SNN (large) 705.01 11.94 0.68
LitE-SNN (medium) 271.41 4.78 0.26
LitE-SNN (small) 129.74 2.39 0.13
@@ 5 Conclusion
In this paper, we propose the LitE-SNN to integrate spatio-temporal compression into the automatic network design. We present the CompConv that expands the search space to support pruning and mixed-precision quantization. Meanwhile, we propose the compressive timestep search to identify the optimal number of timesteps to process SNNs. We also formulate a joint optimization algorithm to learn these compression strategies and architecture parameters. Experimental results show that our LitE-SNN exhibits competitive accuracy, compact model size, and fewer computation costs, making it an attractive choice for deployment on resource-constrained edge devices and opening new possibilities for the practical implementation of SNNs in real-world applications.
Acknowledgements
This work is supported by the Natural Science Foundation of China (No. 61925603, U1909202) and the Key Research and Development Program of Zhejiang Province in China (2020C03004). This work is also supported by IAF, A*STAR, SOITEC, NXP, National University of Singapore under FD-fAbrICS: Joint Lab for FD-SOI Always-on Intelligent & Connected Systems (Award I2001E0053) and the Internal Project Fund from Shenzhen Research Institute of Big Data under Grant T00120220002.